Moriband Corp. paid a dividend of $2.15 yesterday. The company's dividend is expected to grow at a steady rate of 5 percent for the foreseeable future. If investors in stocks of companies like Moriband require a rate of return of 15 percent, what should be the market price of Moriband stock?

X-Centric Energy Company has issued perpetual preferred stock with a stated (par) value of $100 and a dividend of 4.5 percent. If the required rate of return is 8.25 percent, what is the stock's current market price?
Answer #1

1.Current price=D1/(Required return-Growth rate)

=(2.15*1.05)/(0.15-0.05)

=$22.575

2.Annual dividend=100*4.5%=4.5

Hence stock value=Annual dividend/required rate of return

=4.5/0.0825

=$54.55(Approx).